# Distribution-Free Calibration of Statistical Confidence Sets

### v1

@@ 5 Applications
To assess performance, we use several statistic and likelihood/simulator combinations across multiple sample sizes $n$ and simulation budgets $B$ in both likelihood-based and likelihood-free scenarios. In the nuisance parameter setting, we benchmark our methods with two examples: one using likelihood-free inference and another based on likelihood. We set a common confidence level of $0.951-\alpha=$ 0.95 across all experiments. Comparisons are made against the following baseline methods:
@@
We replicate each experiment 50 times for likelihood-based, 30 times for likelihood-free, and 15 times for nuisance examples, and compute the average MAE or deviation from the oracle and its standard error. We identify the top methods in each case as those with the lowest average error, followed by any other methods with performance not significantly different from the best (determined by overlapping 95% asymptotic confidence intervals for the MAE). All computed MAE and standard errors are available in the supplementary material.
@@ 5.1 Examples with tractable likelihood function
In this section, we apply our methods to construct confidence sets on three standard statistical models. We always assume the data is $\mathbf{X}=(X_{1},\ldots,X_{n})$, where the $X_{i}$’s are i.i.d. The models are:
We explore the following choices of $\tau$:
  • Likelihood ratio: Considering that (.$\mathcal{L}(.;\theta_{0})$ represents the likelihood function, the likelihood ratio statistic is given by:
    Under regularity conditions, Wilks’ theorem (Drton, 2009) implies that $-2LR(\mathbf{X},\theta_{0})|\theta=\theta_{0}$ has an asymptotic $\chi_{1}^{2}$ distribution, which is typically used to approximate $C_{\theta}$. However, regularity conditions are not always met, as in the case of the Gaussian Mixture Model (Chen and Li, 2009).
[…]
    and then we compute the Bayes Factor statistic as:
[…]
To compare all methods, we consider $n=1,10,50,100$ and $B=1000,5000,10000,15000$ for all combinations of statistic and likelihood, except for the KS statistic, where we start with $n=5$ to avoid producing a degenerate estimation of the theoretical CDF $F_{\theta_{0}}$. Figure 4 provides a summary of each method’s performance, while Figure 5 shows the coverage control of each method under varying error bounds, considering only the scenarios where asymptotic cutoffs can be derived. The results highlight several key performance differences among the methods:
Figure 4: Frequency of times each method performed best (smallest significant MAE) for each statistic and likelihood model combination across different sample sizes $n$ and simulation sizes $B$. Our methods are highlighted in bold. Overall, our methods demonstrated superior performance compared to other competing approaches in small sample settings, such as $n=1$, $n=5$, and $n=10$, while maintaining competitive performance for larger sample sizes, such as $n=50$ and $n=100$.
@@
  • At $B=1000$ Figure 4 shows no clear advantage among the methods for different sample sizes $n$, with TRUST and tuned TRUST++ showing competitive results in these scenarios. This balance may be attributed to the limited number of simulation samples, which reduces the ability to differentiate among methods regarding cutoff estimation and overall performance.
  • Figure 4 shows that tuned TRUST++ consistently outperforms all competing methods across $B$ from $5000$ to $15000$ and for all $n$. Additionally, Figure 5 highlights the precision and reliability of tuned TRUST++ in achieving conditional coverage. Tuned TRUST++ maintains a high proportion of scenarios with strictly bounded MAE while showcasing lower loss than other competing approaches.
@@
  • TRUST exhibit strong performances for $n\geq 5$, showing a higher count of scenarios with superior outcomes across each combination of $n$ and $B$, as illustrated in Figure 4. Similar to tuned TRUST++, TRUST also exhibits consistency in achieving conditional coverage, as evidenced in Figure 5.
  • Both Monte-Carlo and Asymptotic methods present poor performance when the simulation budget exceeds $1000$ and display an inconsistent control over conditional coverage as evidenced by Figures 4 and 5. Notably, Figure 5 illustrates that the asymptotic method only achieves lower loss values than other approaches in scenarios where it has a low proportion of cases with controlled MAE.
@@
  • Although the boosting method demonstrates competitive coverage consistency and strong performance in certain scenarios and specific combinations of $n$ and $B$, as visualized in Figures 4 and 5, it is clear that this method is generally outperformed by TRUST and tuned TRUST++ in the majority of cases, and also display higher loss in instances where MAE bounds are tighter.
  • Although TRUST++ MV does not match the performance of the tuned TRUST++ and TRUST, it demonstrates strong coverage consistency and performs well in certain scenarios, as shown in Figures 4 and 5.
Figure 5: Top: Frequency of scenarios where each method achieves an MAE below varying fixed thresholds. Notably, as the threshold tightens, our methods continue to exhibit a high count of scenarios with minimal coverage errors. Bottom: Boxplot of MAE for each scenario with error values below the fixed thresholds. As the threshold decreases, both TRUST and TRUST++ demonstrate distinctly lower MAE values compared to competing methods. For comparability, only scenarios permitting asymptotic cutoff derivations were considered.
@@ 5.2 Likelihood-free Inference Problems
We now apply our methods to construct confidence sets across 5 known LFI benchmarks, where the likelihood is intractable or no analytical formula exists for certain statistics, requiring them to be estimated. In cases of intractable likelihood, we simulate the data $\mathbf{X}=(X_{1},\dots,X_{n})$ using a high-fidelity simulator $F_{\theta}$. Given a fixed $\theta$, we also assume that each $X_{i}$ is i.i.d. The simulators and models are:
For choices of $\tau$, we consider both the E-value and BFF introduced in Section 5.1 along with the Waldo statistic (Masserano et al., 2023):
Since an exact posterior $f(\theta|\mathbf{x})$ is unavailable for all statistics in the likelihood-free inference (LFI) context, we estimate posterior quantities using a Neural Posterior Estimator based on normalizing flows (Rezende and Mohamed, 2015; Hermans et al., 2021). This method approximates the posterior $f(\theta|\mathbf{x})$ through an amortized estimator $\hat{f}_{\psi}(\theta|\mathbf{x})$, constructed with neural network-based bijective transformations parameterized by $\psi$ (Rezende and Mohamed, 2015). The estimator is trained on a simulated sample $\{(\theta_{1},\mathbf{X}_{1}),\dots,(\theta_{B},\mathbf{X}_{B})\}$ generated from each model or simulator. Details of normalizing flow architecture, implementation, and simulation budgets are given in Appendix A.
For each statistic and simulator, we evaluate all methods across combinations of $n\in\{1,5,10,20\}$ and $B\in\{10,15,20,30\}\times 10^{3}$, increasing the simulation budgets relative to Section 5.1 to account for the greater complexity of these problems. Figure 6 summarizes the overall performance of each method, while Figure 7 illustrates the coverage control achieved by each method under different error thresholds. The results again reveal several performance distinctions among the methods:
Figure 6: Frequency of times each method performed best for each statistic and simulator combination across different sample sizes $n$ and simulation sizes $B$. Our methods are highlighted in bold. Overall, TRUST++ consistently demonstrated superior performance across all scenarios with a simulation budget exceeding $10000$, while also maintaining competitive results for $B=10000$.
@@
  • Figure 4 demonstrates that tuned TRUST++ consistently outperforms competing methods in nearly all scenarios (15 out of 16). Furthermore, Figure 7 highlights the precision and reliability of tuned TRUST++ in achieving conditional coverage even in likelihood-free scenarios, mirroring its strong performance observed in tractable experiments. Again, tuned TRUST++ consistently exhibits a high count of combinations with tightly bounded MAE and lower overall loss compared to all other approaches.
  • The Monte Carlo method performs poorly across nearly all scenarios, showing weaker coverage control than other methods, as illustrated in Figures 6 and 7. These results highlight the limitations of the Monte Carlo method in adapting to likelihood-free scenarios.
  • The boosting method demonstrates competitive coverage control and consistent performance in several scenarios, notably outperforming in the specific case of $n=1$ and $B=10000$, as shown in Figures 4 and 5. However, in most cases, it is outperformed by both tuned TRUST++ and TRUST++ MV. Aligned with results from tractable experiments, the boosting method also exhibits higher loss under tighter MAE bounds compared to tuned TRUST++ in the likelihood-free context.
    Figure 7: Top: Frequency of LFI scenarios where each method achieves an MAE below varying fixed thresholds. Particularly, as the threshold diminishes, TRUST++ continues to exhibit a high count of scenarios with minimal coverage errors. Bottom: Boxplot of MAE for each scenario with error values below the fixed thresholds. As the threshold decreases, TRUST++ presents a lower MAE distribution compared to all other approaches.
  • Despite exhibiting poorer coverage control than boosting and tuned TRUST++, TRUST++ MV still shows strong performance across all scenarios, particularly standing out in cases with $n=20$. This highlights the effective adaptation of the Majority-Vote scheme for computing local cutoffs within likelihood-free contexts.
  • In contrast to the results seen in tractable experiments, TRUST demonstrates subpar performance and coverage control compared to both boosting and TRUST++ in likelihood-free scenarios, as shown in Figures6 and 7. This indicates that the TRUST partition requires enhancements for likelihood-free contexts, a need effectively addressed through the ensembling approach used in TRUST++.
From these experiments, we conclude that even in challenging likelihood-free examples, TRUST++ excels in achieving conditional coverage.
@@
Figure 8: Visualization of 95% confidence regions and TRUST++ uncertainty for the two moons example, utilizing the e-value statistic with $n=20$ and $B=15000$. In the uncertainty panel, the maroon areas represent regions where TRUST++ exhibits high confidence, while the salmon areas indicate regions with uncertainty. TRUST++ not only closely approximates the oracle set but also captures additional oracle information through its uncertainty regions, offering valuable insights into the precision of the cutoff estimation relative to the simulation budget. The user may choose to increase the number of simulations to decrease the uncertain region.
@@ 5.3.2 Gamma GLM experiment
A standard problem in this setting is to derive a valid confidence interval for a single parameter $\beta_{j}$. A widely used approach is to construct an asymptotic confidence interval based on the marginalized likelihood ratio statistic: $LR(\mathbf{y},\beta_{j})=\log\mathcal{L}(\mathbf{y};\boldsymbol{\widehat{\beta}}_{-j},\beta_{j})-\log\mathcal{L}(\mathbf{y};\boldsymbol{\widehat{\beta}})$, where $\boldsymbol{\widehat{\beta}}$ is the MLE, $\boldsymbol{\widehat{\beta}}_{-j}$ the MLE restricted under a fixed $\beta_{j}$ and (.$\mathcal{L}(.;\boldsymbol{\beta})$ denotes the GLM likelihood as defined in Eq. 18. Using this statistic, an asymptotic confidence interval can be easily derived by referencing the $\chi^{2}_{1}$ distribution to compute $C_{\beta_{j}}$ in a invariant way.
@@ 6 Final Remarks
Furthermore, our methods effectively address inference settings with nuisance parameters, a notable challenge for other approaches, especially in likelihood-free inference. Unlike existing techniques, TRUST and TRUST++ also enable robust uncertainty quantification about the oracle confidence sets, providing valuable insights into whether additional simulated data is needed to achieve more reliable estimates.
Theorem 3 demonstrates that both TRUST and TRUST++ are consistent estimators for quantile regression with respect to any random variable in $\mathbf{x}$. While we applied these methods specifically to quantile-regress the conformal score on $\mathbf{x}$, their utility extends far beyond this particular setting. They can be applied in a variety of contexts, such as constructing predictive intervals in a prediction setting, including for multivariate $Y$. This capability allows for the generation of label-conditional predictive sets that aim to control $\mathbb{P}(Y\in\mathbb{R}(\mathbf{X})\mid Y=y)$, ensuring coverage conditional on the label.
@@ Appendix A Experiment details
For all experiments involving the application of normalizing flows for posterior estimation, we utilized the normflows Python package (Stimper et al., 2023). Specifically, we employed the autoregressive rational quadratic spline architecture for each flow (Durkan et al., 2019). All models were trained on an ACER Nitro 5 AN515-54 using GPU (cuda). The architectural details for both the LFI examples and the nuisance parameter experiment are outlined below.
A.1 LFI posterior estimator
@@
A.2 Poisson Counting Experiment posterior estimator

### v2

@@ 5 Applications
To assess performance, we use several statistic and likelihood/simulator combinations across multiple sample sizes $n$ and simulation budgets $B$ in both likelihood-based and likelihood-free scenarios. In the nuisance parameter setting, we benchmark our methods with two examples: one using likelihood-free inference and another based on likelihood. We set a common confidence level of $1-\alpha=0.95$ across all experiments. Comparisons are made against the following baseline methods:
@@
We replicate each experiment 50 times for likelihood-based, 30 times for likelihood-free, and 15 times for nuisance examples, and compute the average MAE or deviation from the oracle and its standard error. We identify the top methods in each case as those with the lowest average error, followed by any other methods with performance not significantly different from the best (determined by overlapping 95% asymptotic confidence intervals for the MAE). All detailed MAE and standard errors visualizations are available in Appendix A.4.
@@ 5.1 Examples with tractable likelihood function
In this section, we apply our methods to construct confidence sets on three standard statistical models. We always assume the data is $\mathbf{X}=(X_{1},\ldots,X_{n})$, where the $X_{i}$’s are i.i.d. To generate the data, we use the Normal model with fixed variance, the Gaussian mixture model (GMM), and the Lognormal model with both mean and scale as parameters. For the choice of $\tau$, we explore the following statistics: Likelihood ratio (Drton, 2009); Kolmogorov-Smirnov (Marsaglia et al., 2003); Frequentist Bayes Factor (Dalmasso et al., 2021) and E-value (Pereira and Stern, 1999). The detailed setup is described in Appendix A.1.1.
To compare all methods, we consider $n=10,20,50,100$ and $B=1000,5000,10000,15000$ for all combinations of statistic and likelihood. Figure 4 provides a summary of each method’s performance, while Appendix A.4.1 shows a detailed comparison of the models’ performance for each of the statistics in each setting. Figure 4 highlights several key performance differences among the methods:
Figure 4: Tractable Problems: Frequency with which each method achieved the best performance (lowest significant MAE) across all statistic–likelihood combinations and varying $n$ and $B$. Our methods (in bold) consistently outperform competitors, with a clear advantage in smaller-sample settings ($n=10,20$), where asymptotic approaches tend to underperform. For disaggregated results, see Appendix A.4.1.
@@
  • At $B=1000$, there are no clear advantage among the methods for different sample sizes $n$, with TRUST and tuned TRUST++ showing competitive results in these scenarios. This balance may be attributed to the limited number of simulation samples, which reduces the ability to differentiate among methods regarding cutoff estimation and overall performance.
  • Tuned TRUST++ consistently outperforms all competing methods across $B$ from $5000$ to $15000$ and for all $n$.
@@
  • TRUST exhibits good performances for all $n$, showing a higher count of scenarios with superior outcomes across each combination of $n$ and $B$.
  • Both Monte-Carlo and Asymptotic methods present comparatively poor performance when the simulation budget exceeds $1000$.
@@
  • Although the boosting method demonstrates strong performance in certain scenarios and specific combinations of $n$ and $B$, this method is generally outperformed by TRUST and tuned TRUST++ in the majority of cases. The detailed case-by-case analysis in Appendix A.4.1 corroborates these findings.
  • Examining the individual statistics (Appendix A.4.1), our methods show superior performance for the KS statistic, most notably TRUST++ (tuned) and TRUST. For the LR statistic, both the asymptotic approach and our methods achieve satisfactory results. Furthermore, our methods exhibit good performance for BFF and E-value, with boosting representing a competitive alternative in these settings.
@@ 5.2 Likelihood-free Inference Problems
We now apply our methods to construct confidence sets across five well-studied LFI benchmarks, where the likelihood is intractable or no analytical formula exists for certain statistics, requiring them to be estimated. In cases of intractable likelihood, we simulate the data $\mathbf{X}=(X_{1},\dots,X_{n})$ using a high-fidelity simulator $F_{\theta}$. Given a fixed $\theta$, we also assume that each $X_{i}$ is i.i.d. The simulators we use are the following: SLCP (Simple likelihood complex posterior), M/G/1, Weinberg (Hermans et al., 2021); Two Moons (Greenberg et al., 2019); SIR (Lueckmann et al., 2021). We consider the folllowing choices of $\tau$: the E-Value, BFF, and Waldo (Masserano et al., 2023). A detailed description of them can be found in the Appendix A.1.2.
Since an exact posterior $f(\theta|\mathbf{x})$ is unavailable for all statistics in the likelihood-free inference context, we estimate posterior quantities using a Neural Posterior Estimator based on normalizing flows (Rezende and Mohamed, 2015; Hermans et al., 2021). This method approximates the posterior $f(\theta|\mathbf{x})$ through an amortized estimator $\hat{f}_{\psi}(\theta|\mathbf{x})$, constructed with neural network-based bijective transformations parameterized by $\psi$ (Rezende and Mohamed, 2015). The estimator is trained on a simulated sample $\{(\theta_{1},\mathbf{X}_{1}),\dots,(\theta_{B},\mathbf{X}_{B})\}$ generated from each model or simulator. Details of normalizing flow architecture, implementation, and simulation budgets are given in Appendix A.
For each statistic and simulator, we evaluate all methods across combinations of $n\in\{1,5,10,20\}$ and $B\in\{10,15,20,30\}\times 10^{3}$, increasing the simulation budgets relative to Section 5.1 to account for the greater complexity of these problems. Figure 5 presents a summary of the overall performance of each method, while Appendix A.4.2 provides a detailed comparison of coverage results, disaggregated by statistic and data-generating process. The results again reveal several performance distinctions among the methods:
Figure 5: Likelihood-free Inference Problems: Frequency with which each method achieved the best result for each statistic–simulator combination across varying sample sizes $n$ and simulation budgets $B$. Our methods (in bold) show consistent superiority when B>10, [^0] B>10{,}000 and remain competitive even at B=10, [^0] B=10{,}000. For disaggregated results, see Appendix A.4.2.
@@
  • Figure 5 indicates that tuned TRUST++ outperforms competing methods in nearly all combinations (15 out of 16). Again, tuned TRUST++ exhibits a high count of combinations compared to all other approaches, mirroring its performance observed in tractable experiments.
  • The Monte Carlo method performs poorly across nearly all scenarios, showing weaker coverage control than other methods.
  • The boosting method demonstrates competitive coverage control and consistent performance in several scenarios, notably outperforming in the specific case of $n=1$ and $B=10000$. However, in most cases, it is outperformed by both tuned TRUST++ and TRUST++ MV.
  • Despite exhibiting poorer coverage control than boosting and tuned TRUST++, TRUST++ MV still shows good performance across all scenarios.
  • In contrast to the results seen in tractable experiments, TRUST demonstrates subpar performance and coverage control compared to both boosting and TRUST++ in likelihood-free scenarios. This indicates that the TRUST partition requires enhancements for likelihood-free contexts, a need effectively addressed through the ensembling approach used in TRUST++.
  • Appendix A.4.2 further corroborates the effectiveness of our methods, showing that both TRUST++ tuned and TRUST++ MV perform consistently well across all statistics. Boosting also remains a good competitor, particularly for Waldo and E-value. Notably, in most scenarios where either TRUST++ tuned or TRUST++ MV attains the best performance, the other method performs within a narrow margin, effectively placing both among the top-performing approaches.
@@
Figure 6: 95% confidence regions and TRUST++ uncertainty for the two-moons example using the e-value statistic ($n=20$, B=15, [^0] B=15{,}000). Maroon areas denote high confidence, while salmon areas indicate uncertainty. TRUST++ closely matches the oracle set and highlights where additional simulations could reduce uncertainty.
@@ 5.3.2 Gamma GLM experiment
[…]
@@ 6 Final Remarks
Furthermore, our methods effectively address inference settings with nuisance parameters, a notable challenge for other approaches, especially in likelihood-free inference. Unlike existing techniques, TRUST and TRUST++ also enable robust uncertainty quantification about the oracle confidence sets, providing valuable insights into whether additional simulated data is needed to achieve more reliable estimates. Additionally, we can leverage the third branch of LF2I (Dalmasso et al., 2020, 2021) to test for the exact conditional coverage of our confidence sets.
Theorem 3 demonstrates that both TRUST and TRUST++ are consistent estimators for quantile regression. While we applied these methods specifically to quantile-regress the conformal score on $\theta$, their utility extends far beyond this particular setting. They can be applied in a variety of contexts, such as constructing predictive intervals in a prediction setting, including for multivariate $Y$. This capability allows for the generation of label-conditional predictive sets that aim to control $\mathbb{P}(Y\in\mathbb{R}(\mathbf{X})\mid Y=y)$, ensuring coverage conditional on the label.
@@ Appendix A Experiment details
In this section, we provide supplementary details regarding our experimental methodology and present additional results. We specify the statistics and models/simulators used for both the tractable and intractable likelihood comparisons and describe the architectures for the LFI posterior estimator and the estimator used in the Poisson Counting experiment.
For implementation, all experiments involving normalizing flows for posterior estimation were conducted using the normflows Python package (Stimper et al., 2023). We employed an autoregressive rational quadratic spline architecture for each flow (Durkan et al., 2019), and the models were trained on an Acer Nitro 5 AN515-54 using a GPU (CUDA).
A.1 Statistics and models details
A.1.1 Tractable likelihood
The data-generating processes are:
To choose $\tau$, we consider:
  • Likelihood ratio: Considering that $\mathcal{L}(.;\theta_{0})$ represents the likelihood function, the likelihood ratio statistic is given by:
    Under regularity conditions, Wilks’ theorem (Drton, 2009) implies that $2LR(\mathbf{X},\theta_{0})|\theta=\theta_{0}$ has an asymptotic $\chi_{1}^{2}$ distribution, which is typically used to approximate $C_{\theta}$. However, regularity conditions are not always met, as in the case of the Gaussian Mixture Model (Chen and Li, 2009).
    $BF(\mathbf{x},\Theta_{0})=\frac{\mathbb{P}(\mathbf{x}|H_{0})}{\mathbb{P}(\mathbf{x}|H_{1})}=\frac{\int_{\Theta_{0}}\mathcal{L}(\mathbf{x};\theta)d\pi_{0}(\theta)}{\int_{\Theta_{1}}\mathcal{L}(\mathbf{x};\theta)d\pi_{1}(\theta)}\;,$ (15)
    and then we compute the Bayes Factor statistic as:
    $BF(\mathbf{x},\theta_{0})=\frac{\int_{\Theta_{0}}\mathcal{L}(\mathbf{x};\theta)d\pi_{0}(\theta)}{\int_{\Theta_{1}}\mathcal{L}(\mathbf{x};\theta)d\pi_{1}(\theta)}=\frac{\mathcal{L}(\mathbf{x};\theta_{0})}{f(\mathbf{x})}=\frac{f(\theta_{0},\mathbf{x})}{\pi(\theta_{0})\cdot f(\mathbf{x})}=\frac{f(\theta_{0}|\mathbf{x})}{\pi(\theta_{0})}\;.$ (16)
    $ev(\mathbf{x},\theta_{0})=1-\mathbb{P}(\theta\in T_{\theta_{0}}|\mathbf{x})=1-\int_{T_{\theta_{0}}}f(\theta|\mathbf{x})d\theta.$ (17)
A.1.2 Intractable likelihood
We consider the following simulators:
For choices of $\tau$, we consider both the E-value and BFF introduced in Appendix A.1.1 along with the Waldo statistic (Masserano et al., 2023):
A.2 LFI posterior estimator
@@
A.3 Poisson Counting Experiment posterior estimator
@@
A.4 Experiment results
This subsection presents heatmaps comparing the conditional coverage performance of each method for each statistic separately. In each heatmap, rows correspond to the methods and columns represent specific scenarios, defined by a model and budget combination. The heatmaps are divided by each sample size. For each scenario, we highlight the top-performing methods in green, which are those with the lowest MAE or an MAE within two standard errors of the minimum.
A.4.1 Tractable likelihood detailed results
Figures 8, 9, 10, and 11 show the performance of the BFF, LR, KS, and E-value statistics, respectively, for sample sizes of $n=10,20,50,100$.
Overall, TRUST++ tuned and TRUST are strong competitors for tractable models. Nonetheless, certain methods are preferable in specific scenarios. For instance, the asymptotic approach performs particularly well for the LR statistic under the normal model with fixed variance (Figure 9). Similarly, boosting shows competitive performance for the E-value: while not consistently superior across all setups, its performance improves with larger sample sizes. For the E-value, our methods perform well for $n=10,20$, but their advantage diminishes as the sample size increases.
Figure 8: Best methods (lowest MAE) for BFF on tractable models. TRUST++ tuned, TRUST++, and boosting are best performers in general. For $n=10,100$, TRUST is also a decent competitor, and MC is among the best in a few setups, proving to be the least robust choice.
Figure 9: Best methods (lowest MAE) for LR on tractable models. The asymptotic is always among the best competitors for the Normal model with fixed variance D.G.P., and also appears in some setups for the GMM. Again, our methods are consistently good performers, especially TRUST++ tuned and TRUST.
Figure 10: Best methods (lowest MAE) for KS on tractable models. TRUST++ tuned is among the best methods in 47 out of 48 setups, proving to be a powerful competitor. TRUST is also very consistent, appearing in the majority os scenarios. MC and TRUST++ MV have appeared less than the other two, but are considerably more present than boosting and asymptotic, which are not a good choice for this statistic, as shown in the graph.
Figure 11: Best methods (lowest MAE) for E-value on tractable models. In this case, we do not see a specific method that shows a clear dominance as in Figure 10. However, it is possible to notice the poor performance of asymptotic, which is never among the best methods. Our methods and the boosting are consistent across the setups.
A.4.2 Intractable likelihood detailed results
Figures 12, 13, and 14 present the models’ performance for the BFF, E-value, and Waldo statistics, respectively, in the LFI experiments with sample sizes $n=1,5,10,20$.
TRUST++ tuned, TRUST++ MV, and boosting emerge as the overall best performers. TRUST achieves reasonable results in certain scenarios (e.g., Waldo on the SIR benchmark with $n=10$) but rarely ranks among the top methods. MC, in contrast, consistently performs poorly for intractable likelihoods.
For $n=1$, TRUST++ tuned, TRUST++ MV, and boosting achieve strong results across all statistics. Interestingly, TRUST performs well for the E-value in the tractable generator, showing good coverage control for B=15,000,20,000,30, [^0] B=15{,}000,20{,}000,30{,}000.
Figure 12: Best methods (lowest MAE) for BFF on intractable models. TRUST++ tuned, TRUST++ MV and boosting are the strongest competitors in this scenario. In 79 out of 80 experiments, at least one of our methods is highlighted as the best performer. MC and TRUST performances are consistently poorer than their competitors.
Figure 13: Best methods (lowest MAE) for E-value on intractable models. As on BFF, TRUST++ tuned, TRUST++ MV and boosting are the most consistent competitors. However, for this statistic, TRUST stands out in some scenarios, especially when $n=5,10$ (i.e., not in the extreme cases).
Figure 14: Best methods (lowest MAE) for Waldo on intractable models. Similarly to BFF, TRUST++ tuned, TRUST++ MV and boosting are the methods that stand out the most. MC only appears as the winner one time out of 80 experiments, and TRUST has a modest performance, with good results for the SIR benchmark with $n=10$.